7. During an experiment, a signal from a spaceship reached the ground station in five minutes. What was the distance of the spaceship from the ground station? The signal travels at the speed of light, that is, 3 × 108 ms-1.

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Speed = 3 × 108 ms-1 Time = 5 min = 5 × 60 = 300 seconds Distance = Speed × Time Distance = 3 × 108 ms-1 × 300 seconds = 9 × 1010 m